OpenStudy (anonymous):

Use implicit differentiation to find the equation of the tangent line to the curve xy^3+xy= 6 at the point ( 3 , 1 ). The equation of this tangent line can be written in the form y = mx+b where m is: and where b is:

OpenStudy (anonymous):

first you need \(y'\) so find the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

take the derivative using the chain rule and the product rule you should get \[xy^3+xy= 6\] \[y^3+3xy^2y'+y+xy'=0\] ok with that? then we have to solve for \(y'\)

OpenStudy (anonymous):

how do u have the slope?

OpenStudy (anonymous):

we have nothing yet until we solve this equation for \(y'\)

OpenStudy (anonymous):

o ok. i see.

OpenStudy (anonymous):

you have a choice, you can plug in the numbers now for x and y, or you can use algebra to solve for \(y'\) first and then plug in the numbers. either way

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y^3+3xy^2y'+y+xy'=0\] \[3xy^2y'+xy'=-y-y^3\] \[y'(xy^2+x)=-y-y^3\] \[y'=\frac{-y-y^3}{xy^2+x}\] now replace x by 3, y by 1 and find \(y'\) that is your slope. then point- slope formula

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

ok let me know if you get stuck

OpenStudy (anonymous):

slop is -2/9?

OpenStudy (anonymous):

lets see i get \(-\frac{2}{6}=-\frac{1}{3}\) but i could be wrong

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

no -1/3 is no correct as per my hw

OpenStudy (anonymous):

yeah cause i am an idiot and dropped a 3 somewhere you are right

OpenStudy (anonymous):

:) im not totally hopeless in math then .lol

OpenStudy (anonymous):

should be \[y'=\frac{-y-y^3}{3xy^2+x}\]

OpenStudy (anonymous):

should be minus right?

OpenStudy (anonymous):

im not sure. i got -2/9. but my hw still said it was wrong

OpenStudy (anonymous):

lets see if i can get a correct answer \[\frac{-1-1}{3\times 3+3}=\frac{-2}{12}=-\frac{1}{6}\] who about that one?

OpenStudy (anonymous):

1 + 9y' + 1 + 3y' = 0 12y' = -2 ->y' = -2/12 = -16

OpenStudy (anonymous):

ok thats it 1/-6

OpenStudy (anonymous):

I mean -1/6

OpenStudy (anonymous):

@Chlorophyll bless you !

OpenStudy (anonymous):

yes thanks...

OpenStudy (anonymous):

knew we would get there eventually

OpenStudy (anonymous):

@satellite73 I always is the slowest one :")

OpenStudy (anonymous):

yeah and i make all the mistakes!

OpenStudy (anonymous):

3/2 is b

OpenStudy (anonymous):

and i dont know what the hell im doing with math

OpenStudy (anonymous):

now you are using the old point-slope formula \[y-y_1=m(x-x_1)\] with \[x_1=3,y_1=1,m=-\frac{1}{6}\]

OpenStudy (anonymous):

\[y-1=-\frac{1}{6}(x-3)\] etc

OpenStudy (anonymous):

@satellite73 If I spend time to explain like you does, I surely make mistakes at every single line :(